OpenStudy (snowxangel):

Is an equation still solvable if there are variables in the denominator? Isn't there a value for x that will make the denominator zero and then you can't divide by zero?

OpenStudy (snowxangel):

@satellite73

OpenStudy (snowxangel):

@SolomonZelman

OpenStudy (snowxangel):

oh wait forgot you suck at word problems nvm

OpenStudy (anonymous):

lol

OpenStudy (phi):

as long as the solution is not the number that makes the "bottom" zero, you are ok

OpenStudy (anonymous):

iof course they are \[\frac{2}{x}=3\] can be solved right?

OpenStudy (snowxangel):

that's true

OpenStudy (solomonzelman):

if you have something like: \(\normalsize\color{royalblue}{ \frac{\LARGE 3}{\LARGE x}=0 }\) then you won't have solutions.

OpenStudy (snowxangel):

so basically I shouldn't choose a number that will make the denominator zero

OpenStudy (solomonzelman):

But something like: \(\normalsize\color{royalblue}{ \frac{\LARGE 3}{\LARGE x}=4 }\) (or whenever the other side is not a zero, if you have an \(\normalsize\color{royalblue}{ x }\) in the denominator, then there will be a solution)

OpenStudy (snowxangel):

ohhhhhhhhhhhhhhhhhhhhhh

OpenStudy (snowxangel):

ok I think I get it

OpenStudy (snowxangel):

I think

OpenStudy (snowxangel):

yeah kind of

OpenStudy (solomonzelman):

yes, because: \(\normalsize\color{royalblue}{ \frac{\LARGE 3}{\LARGE x}=0 }\) is never true. there is not x value that satisfies this statement

OpenStudy (snowxangel):

ok seems simple now, thanks

OpenStudy (solomonzelman):

\(\normalsize\color{royalblue}{ \frac{\LARGE 3}{\LARGE x}=0 }\) you can multiply both sides times x, and then you get: \(\normalsize\color{royalblue}{ 3=0 }\)

OpenStudy (solomonzelman):

it is either way false

OpenStudy (snowxangel):

ok yeah makes sense

OpenStudy (solomonzelman):

:)